# From Peptides to Nanostructures: A Euclidean Transformer for Fast and Stable Machine Learned Force Fields

### v1

@@ Abstract
Recent years have seen vast progress in the development of machine learned force fields (MLFFs) based on ab-initio reference calculations. Despite achieving low test errors, the suitability of MLFFs in molecular dynamics (MD) simulations is being increasingly scrutinized due to concerns about instability. Our findings suggest a potential connection between MD simulation stability and the presence of equivariant representations in MLFFs, but their computational cost can limit practical advantages they would otherwise bring.
To address this, we propose a transformer architecture called SO3krates that combines sparse equivariant representations (Euclidean variables) with a self-attention mechanism that can separate invariant and equivariant information, eliminating the need for expensive tensor products. SO3krates achieves a unique combination of accuracy, stability, and speed that enables insightful analysis of quantum properties of matter on unprecedented time and system size scales. To showcase this capability, we generate stable MD trajectories for flexible peptides and supra-molecular structures with hundreds of atoms. Furthermore, we investigate the PES topology for medium-sized chainlike molecules (e.g., small peptides) by exploring thousands of minima. Remarkably, SO3krates demonstrates the ability to strike a balance between the conflicting demands of stability and the emergence of new minimum-energy conformations beyond the training data, which is crucial for realistic exploration tasks in the field of biochemistry.
@@ I Introduction
Atomistic modeling relies on long-timescale molecular dynamics (MD) simulations to reveal how experimentally observed macroscopic properties of a system emerge from interactions on the microscopic scale [1]. The predictive accuracy of such simulations is determined by the accuracy of the interatomic forces that drive them. Traditionally, these forces are either obtained from exceedingly approximate mechanistic force fields (FF) or accurate, but computationally prohibitive ab initio electronic structure calculations. Recently, machine learning (ML) potentials have started to bridge this gap, by exploiting statistical dependencies of molecular systems with so far unprecedented flexibility [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@
The accuracy of MLFFs is traditionally determined by their test errors on a few established benchmark datasets [8, 26, 27]. Despite providing an initial estimate of MLFF accuracy, recent research [28, 29, 30] indicates that there is only a weak correlation between MLFF test errors and their performance in long MD simulations, which is considered the true measure of predictive usefulness. Faithful representations of dynamical and thermodynamic observables can only be derived from accurate MD trajectories. From an ML perspective this shortcoming can be attributed to poor extrapolation behavior, which becomes particularly severe for high temperature configurations or conformationally flexible structures. In these cases, the geometries explored during MD simulations significantly deviate from the distribution of the training data.
@@
Figure 1: (a) Illustration of an invariant convolution, an SO(3) convolution and of the Euclidean attention mechanism that underlies the SO3krates transformer. We decompose the representation of molecular structure into high dimensional invariant features and equivariant Euclidean variables (EV), which interact via self-attention. (b) The proposed design paradigm can help to overcome current trade-offs between stability in MD simulations and computational efficiency experienced for other (equivariant) MPNNs. (c) Computational efficiency of SO3krates allows the calculation of velocity-auto correlation functions from converged MD simulations for supra-molecular structures. (d) SO3krates enables to explore thousands of minima of the potential energy surface of small chainlike molecules such as Ac-Ala3-NHMe or DHA, where SO3krates can robustly extrapolate beyond the training data.
@@
The ongoing progress in MLFF development has resulted in a wide range of increasingly sophisticated model architectures aiming to improve the extrapolation behavior. Among these, message passing neural networks (MPNNs) [9, 31, 12] have emerged as a particularly effective class of architectures. MPNNs can be considered as a generalization of convolutions to handle unstructured data domains, such as molecular graphs. This operation provides an effective way to extract features from the input data and is ubiquitous in many modern ML architectures. Recent advances in this area focused on the incorporation of physically meaningful geometric priors [8, 11, 32, 23, 33]. This has lead to so-called equivariant MPNNs, which have been found to reduce the obtained approximation error [34, 33, 35, 36] and offer better data efficiencies than invariant models [33]. Invariant models rely on pairwise distances to describe atomic interactions, as they do not change upon rotation [5]. However, with growing system size, flexibility or chemical heterogeneity, it becomes increasingly harder to derive the correct interaction patterns within this limited representation. This is why equivariant models enable to incorporate additional directional information, to capture interactions depending on the relative orientation of neighboring atoms. It allows them to discriminate interactions that can appear inseparable to simpler models [34] and to learn more transferable interaction patterns from the same training data.
@@
SchNet [9] $\mathcal{O}(n\times\braket{\mathcal{N}}\times F)$ 0
PaiNN [34] $\mathcal{O}(n\times\braket{\mathcal{N}}\times l_{\mathrm{max}}^{2}\times F)$ 1
SpookyNet [24] $\mathcal{O}(n\times\braket{\mathcal{N}}\times l_{\mathrm{max}}^{2}\times F)$ 2
NequIP [33] $\mathcal{O}(n\times\braket{\mathcal{N}}\times l_{\mathrm{max}}^{6}\times F)$ 3
SO3krates $\mathcal{O}(n\times\braket{\mathcal{N}}\times(l_{\mathrm{max}}^{2}+F))$ 3
@@
Table 1: Scaling for different (equivariant) message passing architectures, where $n$ is the number of atoms, $\braket{\mathcal{N}}$ the average number of neighbors and $l_{\mathrm{max}}$ the maximal degree.
@@
We take this as motivation to propose an Euclidean self-attention mechanism that replaces SO(3) convolutions with a filter on the relative orientation of atomic neighborhoods, representing atomic interactions without the need for expensive tensor products. Our solution builds on recent advances in neural network architecture design [39] and from the field of geometric deep learning [40, 34, 33, 35]. Our SO3krates method uses a sparse representation for the molecular geometry and restricts projections of all convolution responses to the most relevant invariant component of the equivariant basis functions. Due to the orthonormality of the spherical harmonics, such a projection corresponds to partial traces of the product-tensor, which can be expressed in terms of linear-scaling inner products. This enables efficient scaling to high degree equivariant representations without sacrificing computational speed and memory cost. Force predictions are obtained from the gradient of the resulting invariant energy model, which represents a piece-wise linearization that is naturally equivariant. Throughout, a self-attention mechanism is used to decouple invariant and equivariant basis elements within the model.
@@
We compare the stability and speed of the proposed SO3krates model with current state-of-the art ML potentials and find that our solution overcomes the limitations of current equivariant MLFFs, without compromising on their advantages. Our proposed mathematical formulation leading to an efficient equivariant architecture enables reliably stable MD simulations with a speedup of up to a factor of $\sim 25$ over equivariant MPNNs with comparable stability and accuracy [30].
@@
To demonstrate this, we run accurate nanosecond-long MD simulations for supra-molecular structures within only a few hours, which allows us to calculate converged auto-correlation functions (vibrational spectra) for structures that range from small peptides with 42 atoms up to nanostructures with 370 atoms. We further apply our model to explore the topology of the PES of docosahexaenoic acid (DHA) and Ac-Ala3-NHMe by investigating 10k minima using a minima hopping algorithm [41]. Such an investigation requires roughly 30M FF evaluations that are queried at temperatures between a few 100 K up to $\sim\!1200$ K. With DFT methods, this analysis would require more than a year of computation time. Existing equivariant MLFFs with comparable prediction accuracy would run more than a month for such an analysis. In contrast, we are able to perform the simulation in only 2.5 days, opening up the possibility to explore hundreds of thousands of PES minima on practical timescales. In one of our experiments, we further show that SO3krates enables the detection of physically valid minima conformations which have not been part of the training data. The ability to extrapolate to unknown parts of the PES is essential for scaling MLFFs to large structures, since the availability of ab-initio reference data can only cover sub-regions for conformationally rich structures.
@@
Furthermore, we examine the impact of disabling the equivariance property in our network architecture to gain a deeper understanding of its influence on the characteristics of the model and its reliability in MD simulations. We find, that the equivariant nature can be linked to the stability of the resulting MD simulation and to the extrapolation behavior to higher temperatures. We are able to show, that equivariance lowers the spread in the error distribution even when the test error estimate is the same on average. Thus, using directional information via equivariant representations shows analogies in spirit to classical ML theory, where mapping into higher dimensions yields richer features spaces that are easier to parametrize [42, 43, 44].
@@ II Results
One way of incorporating the rotational invariance of the PES is to build messages that are based on invariant inputs such as distances, angles or dihedral angles. However, this incomplete list of features can not discriminate certain interaction patterns [48]. An alternative is to use SO(3) equivariant representations [49, 33, 47, 35] within a basis that allows for systematic expansion to match the complexity of the modelled system.
@@
$m_{ij}^{LM}=\sum_{l_{1}l_{2}m_{1}m_{2}}C_{l_{1}l_{2}L}^{m_{1}m_{2}M}\phi^{l_{1}l_{2}L}(r_{ij})Y_{l_{1}}^{m_{1}}(\hat{r}_{ij})f_{j}^{l2m2},$ (3)
where $C_{l_{1}l_{2}L}^{m_{1}m_{2}M}$ are the Clebsch-Gordan coefficients, $Y^{l}_{m}$ is a spherical harmonic of degree $l$ and order $m$, the function $\phi^{l_{1}l_{2}}:\mathbb{R}\mapsto\mathbb{R}^{F}$ modulates the radial part and $f_{j}^{l_{2}m_{2}}\in\mathbb{R}^{F}$ is an atomic feature vector. Thus, performing a single convolution scales as $\mathcal{O}(l_{\mathrm{max}}^{6}\times F)$, where $l_{\mathrm{max}}$ is the largest degree in the network (Fig. 2).
Here we propose two conceptual changes to Eq. (3) that we will denote as Euclidean self-attention: (1) We separate the message into an invariant and an equivariant part and (2) replace the SO(3) convolution by and attention function on its invariant output. To do so, we start by initializing atomic features $f_{i}^{[t=0]}\in\mathbb{R}^{F}$ and Euclidean variables (EV) $x_{i,LM}^{[t=0]}\in\mathbb{R}$ from the atomic types and the atomic neighborhoods, respectively. Collecting all orders and degrees for the EV in a single vector, gives $(l_{\mathrm{max}}+1)^{2}$ dimensional representations $\bm{x}_{i}$ that transform equivariant under rotation and capture directional information up to degree $l_{\mathrm{max}}$ (methods section IV.1).
(1) The message for the invariant part is written as
$m_{ij}=\alpha_{ij}f_{j}$ (4)
and the one for the equivariant part as
@@
where $\alpha_{ij}\in\mathbb{R}$ are (per-degree) attention coefficients. Features and EV are updated with the aggregated messages, which writes as
$f_{i}^{[t+1]}=f_{i}^{[t]}+\sum_{j\in\mathcal{N}(i)}m_{ij}$ (6)
for the features and as
$x_{iLM}^{[t+1]}=x_{iLM}^{[t]}+\sum_{j\in\mathcal{N}(i)}m_{ijLM}.$ (7)
for the EV. Due to the separation, the overall message calculation scales as $\mathcal{O}(l_{\mathrm{max}}^{2}+F)$, replacing the multiplication of feature dimension and $l_{\mathrm{max}}$ from other equivariant architectures by addition (Tab. 1).
@@
where each Euclidean transformer block (ecTblock) consists of a self-attention block and an interaction block. The self-attention block, implements the Euclidean self-attention mechanism described in the former section. The interaction block gives additional freedom for parametrization by exchanging information between features and EV located at the same atom. After $T$ MP steps, per-atom energies $E_{i}$ are calculated from the final features $f_{i}^{[T]}$ using a two-layered neural network and are summed to the total potential energy (Eq. (1)). Atomic forces are obtained using automatic differentiation, which ensures energy conservation. A detailed outline of the architectural components and the proposed Euclidean self-attention framework is given in the methods section.
@@ II.2 Overcoming Accuracy-Stability-Speed Trade-Offs
Figure 4: (a) Number of frames per second (FPS) vs. the averaged stability coefficient (Eq. (28)) in MD simulations run with different state-of-the-art MPNN architectures [33, 12, 34, 50, 38, 45, 51, 52] and (b) FPS vs. the averaged force MAE for four small organic molecules from the MD17 data set as reported in [30]. SO3krates yields reliable MD simulations and high accuracies without sacrificing computational performance. (c) Stability and speed of SO3krates enable nanosecond long MD simulations for supra molecular structures within a few hours. For the buckyball catcher, the ball stays in the catcher over the full simulation time of 20 ns, illustrating that the model successfully picks up on weak, non-covalent bonding.
@@
We show in the following experiment, that SO3krates can overcome the trade-offs between MD stability, accuracy and computational efficiency (Fig. 4).
@@
A recent study compared the stability of different state-of-the-art MLFFs in short MD simulations and found that only the SO(3) convolution based architecture NequIP [33] gave reliably stable results [30]. The excellent stability of such models, however, comes at the price of extensive computational cost (Fig. 4 (a)) which stems from equivariant features and SO(3) convolutions. This leads to a trade-off between the stability and the computational efficiency of MP based MLFFs, but SO3krates can overcome this stability-speed trade-off. It allows to predict up to one order of magnitude more frames per second (FPS) without sacrificing reliability in MD simulations. Although the test accuracy does not necessarily correlate with the stability (compare e.g. GemNet and SphereNet in Fig. 4 (a) and (b)), only accurate and stable models are of ultimate interest. We find, that SO3krates yields accurate force predictions, thus overcoming the complementary trade-off between accuracy and speed (Fig. 4 (b)).
@@
For the radial distribution functions (RDFs), we find consistent results across five simulations (SI Fig. 12) for all of the four investigated structures, which are in agreement with the RDFs from DFT calculations. Interestingly, it has been found that other approaches with a larger number of FPS can give inaccurate RDFs, which result in MAEs between 0.35 for salicylic acid and 1.02 for naphthalene [30]. In comparison the achieved accuracies with SO3krates show that the seemingly contradictory requirements of high computational speed and accurate observables from MD trajectories can be reconciled.

### v2

@@ Abstract
Recent years have seen vast progress in the development of machine learned force fields (MLFFs) based on ab-initio reference calculations. Despite achieving low test errors, the reliability of MLFFs in molecular dynamics (MD) simulations is facing growing scrutiny due to concerns about instability over extended simulation timescales. Our findings suggest a potential connection between robustness to cumulative inaccuracies and the use of equivariant representations in MLFFs, but the computational cost associated with these representations can limit this advantage in practice.
To address this, we propose a transformer architecture called SO3krates that combines sparse equivariant representations (Euclidean variables) with a self-attention mechanism that separates invariant and equivariant information, eliminating the need for expensive tensor products. SO3krates achieves a unique combination of accuracy, stability, and speed that enables insightful analysis of quantum properties of matter on extended time and system size scales. To showcase this capability, we generate stable MD trajectories for flexible peptides and supra-molecular structures with hundreds of atoms. Furthermore, we investigate the PES topology for medium-sized chainlike molecules (e.g., small peptides) by exploring thousands of minima. Remarkably, SO3krates demonstrates the ability to strike a balance between the conflicting demands of stability and the emergence of new minimum-energy conformations beyond the training data, which is crucial for realistic exploration tasks in the field of biochemistry.
@@ I Introduction
Atomistic modeling relies on long-timescale molecular dynamics (MD) simulations to reveal how experimentally observed macroscopic properties of a system emerge from interactions on the microscopic scale [1]. The predictive accuracy of such simulations is determined by the accuracy of the interatomic forces that drive them. Traditionally, these forces are either obtained from exceedingly approximate mechanistic force fields (FF) or accurate, but computationally prohibitive ab initio electronic structure calculations. Recently, machine learning (ML) potentials have started to bridge this gap, by exploiting statistical dependencies of molecular systems with so far unprecedented flexibility [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
The accuracy of MLFFs is traditionally determined by their test errors on a few established benchmark datasets [8, 28, 29]. Despite providing an initial estimate of MLFF accuracy, recent research [30, 31, 32, 33] indicates that there is only a weak correlation between MLFF test errors and their performance in long-timescale MD simulations, which is considered the true measure of predictive usefulness. Faithful representations of dynamical and thermodynamic observables can only be derived from accurate MD trajectories. From an ML perspective this shortcoming can be attributed to poor extrapolation behavior, which becomes particularly severe for high temperature configurations or conformationally flexible structures. In these cases, the geometries explored during MD simulations significantly deviate from the distribution of the training data.
@@
Figure 1: (a) Illustration of an invariant convolution, an SO(3) convolution and of the Euclidean attention mechanism that underlies the SO3krates transformer. We decompose the representation of molecular structure into high dimensional invariant features and equivariant Euclidean variables (EV), which interact via self-attention. (b) Computational efficiency of SO3krates allows the calculation of velocity-auto correlation functions from converged MD simulations for supra-molecular structures. (c) SO3krates enables to explore thousands of minima of the potential energy surface of small chainlike molecules such as Ac-Ala3-NHMe or DHA, where SO3krates can robustly extrapolate beyond the training data.
@@
The ongoing progress in MLFF development has resulted in a wide range of increasingly sophisticated model architectures aiming to improve the extrapolation behavior. Among these, message passing neural networks (MPNNs) [9, 34, 12] have emerged as a particularly effective class of architectures. MPNNs can be considered as a generalization of convolutions to handle unstructured data domains, such as molecular graphs. This operation provides an effective way to extract features from the input data and is ubiquitous in many modern ML architectures. Recent advances in this area focused on the incorporation of physically meaningful geometric priors [8, 11, 35, 23, 36]. This has lead to so-called equivariant MPNNs, which have been found to reduce the obtained approximation error [37, 38, 36, 39, 40] and offer better data efficiencies than invariant models [36]. Invariant models rely on pairwise distances to describe atomic interactions, as they do not change upon rotation [5]. However, with growing system size, flexibility or chemical heterogeneity, it becomes increasingly harder to derive the correct interaction patterns within this limited representation. This is why equivariant models enable to incorporate additional directional information, to capture interactions depending on the relative orientation of neighboring atoms. It allows them to discriminate interactions that can appear inseparable to simpler models [37, 41] and to learn more transferable interaction patterns from the same training data.
@@
SchNet [9] $\mathcal{O}(n\times\braket{\mathcal{N}}\times{\color[rgb]{0,0,0}1}\times F)$ 0
PaiNN [37] $\mathcal{O}(n\times\braket{\mathcal{N}}\times{\color[rgb]{0,0,0}4}\times F)$ 1
SpookyNet [24] $\mathcal{O}(n\times\braket{\mathcal{N}}\times{\color[rgb]{0,0,0}(l_{\mathrm{max}}+1)^{2}}\times F)$ 2
NequIP [36] $\mathcal{O}(n\times\braket{\mathcal{N}}\times{\color[rgb]{0,0,0}(l_{\mathrm{max}}+1)^{6}}\times F)$ 3
SO3krates $\mathcal{O}(n\times\braket{\mathcal{N}}\times({\color[rgb]{0,0,0}(l_{\mathrm{max}}+1)^{2}}+F))$ 3
@@
Table 1: Scaling for different message passing architectures, where $n$ is the number of atoms, $\braket{\mathcal{N}}$ the average number of neighbors and $l_{\mathrm{max}}$ the maximal degree. SchNet andPaiNN have fixed maximal degree of $l_{\mathrm{max}}=0$ and $l_{\mathrm{max}}=1$ whereas they are free parameter in other models.
@@
We take this as motivation to propose an Euclidean self-attention mechanism that replaces SO(3) convolutions with a filter on the relative orientation of atomic neighborhoods, representing atomic interactions without the need for expensive tensor products. Our solution builds on recent advances in neural network architecture design [44] and from the field of geometric deep learning [45, 37, 36, 39]. Our SO3krates method uses a sparse representation for the molecular geometry and restricts projections of all convolution responses to the most relevant invariant component of the equivariant basis functions. Due to the orthonormality of the spherical harmonics, such a projection corresponds to partial traces of the product-tensor, which can be expressed in terms of linear-scaling inner products. This enables efficient scaling to high-degree equivariant representations without sacrificing computational speed and memory cost. Force predictions are obtained from the gradient of the resulting invariant energy model, which represents a piece-wise linearization that is naturally equivariant. Throughout, a self-attention mechanism is used to decouple invariant and equivariant basis elements within the model.
@@
We compare the stability and speed of the proposed SO3krates model with current state-of-the art ML potentials and find that our solution overcomes the limitations of current equivariant MLFFs, without compromising on their advantages. Our proposed mathematical formulation leading to an efficient equivariant architecture enables reliably stable MD simulations with a speedup of up to a factor of $\sim 30$ (Fig. 4 (c)) over equivariant MPNNs with comparable stability and accuracy [32].
@@
To demonstrate this, we run accurate nanosecond-long MD simulations for supra-molecular structures within only a few hours, which allows us to calculate Fourier transforms of converged velocity auto-correlation functions for structures that range from small peptides with 42 atoms up to nanostructures with 370 atoms. We further apply our model to explore the topology of the PES of docosahexaenoic acid (DHA) and Ac-Ala3-NHMe by investigating 10k minima using a minima hopping algorithm [46]. Such an investigation requires roughly 30M FF evaluations that are queried at temperatures between a few 100 K up to $\sim\!1200$ K. With DFT methods, this analysis would require more than a year of computation time. Existing equivariant MLFFs with comparable prediction accuracy would run more than a month for such an analysis. In contrast, we are able to perform the simulation in only 2.5 days, opening up the possibility to explore hundreds of thousands of PES minima on practical timescales. Furthermore, we show that SO3krates enables the detection of physically valid minima conformations which have not been part of the training data. The ability to extrapolate to unknown parts of the PES is essential for scaling MLFFs to large structures, since the availability of ab-initio reference data can only cover sub-regions for conformationally rich structures.
@@
We also examine the impact of disabling the equivariance property in our network architecture to gain a deeper understanding of its influence on the characteristics of the model and its reliability in MD simulations. Here we find, that the equivariant nature can be linked to the stability of the resulting MD simulation and to the extrapolation behavior to higher temperatures. We are able to show, that equivariance lowers the spread in the error distribution even when the test error estimate is the same on average. Thus, using directional information via equivariant representations shows analogies in spirit to classical ML theory, where mapping into higher dimensions yields richer features spaces that are easier to parametrize [47, 48, 49].
@@ II Results
One way of incorporating the rotational invariance of the PES is to build messages that are based on potentially incomplete sets of invariant inputs such as distances, angles or dihedral angles. An alternative is to use SO(3) equivariant representations [38, 36, 52, 39] within a basis that allows for systematic multipole expansion of the geometry to match the complexity of the modelled system.
@@
$m_{ij}^{LM}=\sum_{l_{1}l_{2}m_{1}m_{2}}C_{l_{1}l_{2}L}^{m_{1}m_{2}M}\phi^{l_{1}l_{2}L}(r_{ij})Y_{l_{1}}^{m_{1}}(\hat{r}_{ij})f_{j}^{l_{2}m_{2}},$ (3)
where $C_{l_{1}l_{2}L}^{m_{1}m_{2}M}$ are the Clebsch-Gordan coefficients, $Y^{l}_{m}$ is a spherical harmonic of degree $l$ and order $m$, the function $\phi^{l_{1}l_{2}L}:\mathbb{R}\mapsto\mathbb{R}^{F}$ modulates the radial part and $f_{j}^{l_{2}m_{2}}\in\mathbb{R}^{F}$ is an atomic feature vector. Thus, performing a single convolution scales as $\mathcal{O}(l_{\mathrm{max}}^{6}\times F)$, where $l_{\mathrm{max}}$ is the largest degree in the network (Fig. 2). Here we propose two conceptual changes to Eq. (3) that we will denote as Euclidean self-attention: (1) We separate the message into an invariant and an equivariant part and (2) replace the SO(3) convolution by an attention function on its invariant output. To do so, we start by initializing atomic features $f_{i}^{[t=0]}\in\mathbb{R}^{F}$ and Euclidean variables (EV) $x_{i,LM}^{[t=0]}\in\mathbb{R}$ from the atomic types and the atomic neighborhoods, respectively. Collecting all orders and degrees for the EV in a single vector, gives $(l_{\mathrm{max}}+1)^{2}$ dimensional representations $\bm{x}_{i}$ that transform equivariant under rotation and capture directional information up to degree $l_{\mathrm{max}}$ (methods section IV.1).
(1) The message for the invariant components is expressed as
$m_{ij}=\alpha_{ij}f_{j},$ (4)
whereas the equivariant parts propagate as
@@
where $\alpha_{ij}\in\mathbb{R}$ are (per-degree) attention coefficients. Features and EV are updated with the aggregated messages to
$f_{i}^{[t+1]}=f_{i}^{[t]}+\sum_{j\in\mathcal{N}(i)}m_{ij},$ (6)
for the features and
$x_{iLM}^{[t+1]}=x_{iLM}^{[t]}+\sum_{j\in\mathcal{N}(i)}m_{ijLM},$ (7)
respectively. Due to this separation, the overall message calculation scales as $\mathcal{O}(l_{\mathrm{max}}^{2}+F)$, as it replaces the multiplication of feature dimension and $l_{\mathrm{max}}$ that appears in other equivariant architectures by an addition (Tab. 1). As shown in [53], a separation between invariant and equivariant representations can also achieved by adding an invariant latent space that is updated using iterated tensor products on an equivariant, edge based feature space. In contrast, our approach is centered around atom-wise representations and the a priori separation of both interaction spaces allows to fully avoid the usage of tensor products. Both design choices benefit computational efficiency.
@@
where each Euclidean transformer block (ecTblock) consists of a self-attention block and an interaction block. The self-attention block, implements the Euclidean self-attention mechanism described in the previous section. The interaction block gives additional freedom for parametrization by exchanging information between features and EV located at the same atom. After $T$ MP steps, per-atom energies $E_{i}$ are calculated from the final features $f_{i}^{[T]}$ using a two-layered neural network and are summed to the total potential energy (Eq. (1)). Atomic forces are obtained using automatic differentiation, which ensures energy conservation [8].
We remark that the outlined equivariant architecture does not preclude the modelling of vectorial and tensorial properties, such as atomic quadrupoles or octopoles, up to the set maximum degree $l_{\mathrm{max}}$. For example, molecular dipoles can be learned by combining invariant partial charge predictions with atomic dipoles extracted from the EVs of degree $l=1$ [24, 37].
A detailed outline of the architectural components and the proposed Euclidean self-attention framework is given in the methods section.
@@ II.2 Overcoming Accuracy-Stability-Speed Trade-Offs
Figure 4: (a) Number of frames per second (FPS) vs. the averaged stability coefficient (upper panel) and FPS vs. the averaged force MAE (lower panel) for four small organic molecules from the MD17 data set as reported in [32] for different state-of-the-art MPNN architectures [36, 12, 37, 54, 43, 50, 55, 56]. (b) Since run times are sensitive to hardware and software implementation details, we re-implement two representative models along the trade-off lines under settings identical to the SO3krates MLFF (using jax), which yields framework-corrected FPS (dahshed vs. solid line). We observe speed-ups between 28 (for NequIP) and 15 (for SchNet) in our re-implementations. We find, that SO3krates enables reliable MD simulations and high accuracies without sacrificing computational performance. (c) MD step time vs. the number of atoms in the system. The smaller pre-factor in the computational complexity compared to SO(3) convolutions (Tab. 1) results in computational speed-ups that grow in system size. (d) Stability and speed of SO3krates enable nanosecond-long MD simulations for supra-molecular structures within a few hours. For the buckyball catcher, the ball stays in the catcher over the full simulation time of 20 ns, illustrating that the model successfully picks up on weak, non-covalent bonding.
@@
We now demonstrate in the following numerical experiment, that SO3krates can overcome the trade-offs between MD stability, accuracy and computational efficiency (Fig. 4).
@@
A recent study compared the stability of different state-of-the-art MLFFs in short MD simulations and found that only the SO(3) convolution based architecture NequIP [36] gave reliable results [32]. However, the excellent stability of such models comes at a substantial computational cost associated with this operation (Fig. 4 (a), top panel). This necessitates a trade-off between the stability and the computational efficiency of the MLFF, which SO3krates can now overcome (Fig. 4 (b)). Our model allows the prediction of up to one order of magnitude more frames per second (FPS) (Fig. 4 (c)), enabling step times at sub-millisecond speed, without sacrificing reliability or accuracy in MD simulations (Fig. 4 (b)). We remark however, that test accuracy and stability do not necessarily correlate (compare e.g. GemNet and SphereNet in Fig. 4 (a)) but only simultanously accurate and stable models are of practical interest. We find, that SO3krates yields accurate force predictions, thus overcoming this trade-off effectively (Fig. 4 (b)). As for stability and speed, the investigated models in [32] show an accuracy-speed trade-off (Fig. 4 (a) lower panel) in line with the findings reported in [57].
Any empirical runtime measurement depends on specific hardware and software conditions. The run times reported in [32] have been measured for MLFF models implemented in PyTorch+ASE. To ensure comparability with SO3krates  (implemented in jax), we re-implement two representative models along the trade-off lines (Fig. 4 (a)) under SO3krates-identical settings. Since the study mentioned above reports that the fastest model (ForceNet) yields wrong observables in their benchmark, we instead chose the second fastest contender (SchNet) for our re-implementation. As the most stable and accurate model we chose NequIP for re-implementation. This selection of architectures is also representative for invariant (SchNet) and equivariant SO(3)-convolution-based models (NequIP), constituting the upper and lower bounds in terms of computational complexity (Tab. 1). All models are re-implemented in jax [58] using the e3x library [59]. MD step times are measured with the same MD code written in jax-md [60] on the same physical device (Nvidia V100 GPU). The models follow the default MD17 hyperparameters as outlined in the original publications [12, 36]. This ensures equal footing for our runtime comparisons that follow. Interestingly, the transition from pyTorch+ase (Fig. 4 (b) purple dashed line) to jax+jax-md (Fig. 4 (b) purple solid line) allows for a speed-up between 28 for NequIP and 15 for SchNet. This illustrates the importance of identical settings and the potential of the jax ecosystem. Notably, the step times are measured without the time required for IO operations, since they are highly depended on the local HPC infrastructure. Thus, wall times we report for full simulations have a constant offset w. r. t. the reported step times.
For small organic molecules with up to 21 atoms SO3krates achieves an averaged speed-up by a factor of 5 compared to the NequIP architecture whereas step times are slightly larger (by a factor of 1.4) than for the invariant SchNet model. The speed-up over SO(3) convolutions increases in the total number of atoms (Fig. 4 (c)), which is in line with the smaller pre-factor in the theoretical scaling analysis (Tab. 1) such that for the double walled nanotube (370 atoms), the speed-up compared to  NequIP has grown to a factor of 30. Compared to invariant convolutions, we find our approach to yield slightly slower prediction speed which is in line with theoretical considerations.
@@
For the radial distribution functions (RDFs), we find consistent results across five simulation runs (SI Fig. 14) for all of the four investigated structures, which are in agreement with the RDFs from DFT calculations. Interestingly, it has been found that some faster MLFF models can give inaccurate RDFs, which result in MAEs between 0.35 for salicylic acid and 1.02 for naphthalene [32]. In comparison the achieved accuracies with SO3krates show that the seemingly contradictory requirements of high computational speed and accurate observables from MD trajectories can be reconciled.